OpenStudy (anonymous):

(2.5 x 10^-10) x (1.2 x 10^18)

OpenStudy (anonymous):

300000000

OpenStudy (lgbasallote):

you can rewrite it as \[\huge (2.5 \times 1.2) \times (10^{-10} \times 10^{18})\] is that easier?

OpenStudy (anonymous):

Yes! I was wondering if I could combine those... thanks so much!

OpenStudy (lgbasallote):

yes. it's the commutative/associative property of multiplication